From a batch containing 4 boys and 6 girls, a group of 4 students is to be selected. There are how many combinations of ⦁ 2 boys and 2 girls. ⦁ 3 boys and 1 girl. ⦁ 4 boys. ⦁ At least 2 boys Assignment expert


1
Expert's answer
2021-06-01T03:45:40-0400

1.


"\dbinom{4}{2}\dbinom{6}{2}=6(15)=90"

2.


"\dbinom{4}{3}\dbinom{6}{1}=4(6)=24"

3.


"\dbinom{4}{4}\dbinom{6}{0}=1(1)=1"

4.


"\dbinom{4}{2}\dbinom{6}{2}+\dbinom{4}{3}\dbinom{6}{1}+\dbinom{4}{4}\dbinom{6}{0}"

"=90+24+1=115"